# Lessons Learned from Developing a Privacy-Preserving Multimodal Wearable for Local Voice-and-Vision Inference

### v1

@@ 3.6. Training and Profiling
To design a model that fits the constraints of our ear-mounted device, we trained and profiled it using a commercial TinyML platform, Edge Impulse, targeting reference microcontrollers comparable to the Xiao ESP32-S3 Sense used in our prototype. The wake word detector operates on 1 s windows of 16 kHz mono audio with a 500 ms stride, producing overlapping predictions suitable for continuous listening.
@@
The quantized model achieved 73.9 % validation accuracy (AUC = 0.92), with the strongest separation between the wake word and background noise. Remaining confusions occurred mainly between “hey dotty” and near-rhymes such as “hey dobby,” which we treated conservatively as non-triggers. Despite the synthetic nature of the dataset, this performance was sufficient for robust real-time triggering on-device.
@@ 4. DISCUSSION
Responsiveness on a constrained form factor (RQ1). Achieving interactive latency within a small ear-mounted device required careful partitioning. By restricting the wearable to wake word detection, capture, and streaming—and offloading all inference to a paired smartphone—we kept the device lightweight and thermally safe. Despite limited battery capacity and compute, we achieved query latencies of 2–3 s, showing that fully local multimodal interaction is feasible without cloud support.
Balancing energy, latency, and bandwidth (RQ2). Our design adopts a hybrid sensing pipeline: BLE for provisioning, Wi-Fi for runtime streaming, and short-lived event-driven sessions to reduce idle drain. Audio is compressed via 4-bit ADPCM, and single-frame snapshots replace continuous video. These decisions reduced power and network load while maintaining interactive performance—critical for short-session usage on a 200 mAh battery.
Designing for privacy and acceptability (RQ3). Three properties shaped trust: (i) interaction is explicitly triggered by the user via wake word, (ii) all sensor data is processed locally, with no cloud calls, and (iii) the form factor is subtle and familiar—more pen than camera. This design made the device feel unobtrusive in public. Still, some ambiguity remained; future iterations may include subtle LED or haptic cues to clarify capture status for both wearer and bystanders.
@@
System co-design. Many integration challenges (e.g., Wi-Fi/BLE coexistence, buffer stability, antenna layout) only emerged late in iteration. Prototyping tools like the early “pen webserver” proved invaluable, letting us validate hardware, streaming paths, and thermal behavior before inference was integrated.
@@ 4.1. Limitations
Despite these advances, several limitations remain. The evaluation involved a small number of short-duration deployments and relied primarily on internal testing. Longer, in-the-wild studies are necessary to assess reliability, comfort, and social perception over time. Our analysis also focused on a single smartphone class; behavior on lower-end hardware or across operating systems may differ. Finally, while the current prototype supports wake word activation and multimodal inference, it does not yet integrate higher-level capabilities such as contextual memory or continuous dialogue—areas we plan to address next.
@@ 4.2. Future Work
We are in the process of obtaining IRB approval for a pilot study to evaluate everyday usability. We plan to collect both quantitative and qualitative data, including end-to-end latency, wake word accuracy per-participant, Wi-Fi reconnection behavior, battery drop over sessions, and participant feedback on comfort, perceived privacy, and bystander reactions. Future work will expand both the scope and depth of the system. On the hardware and networking side, we plan extended deployments, adaptive power management, and modular enclosures to support additional sensors and continuous use. Alternatives to wake word activation will also be explored, including integration of a tactile sensor. On the intelligence layer, we aim to develop additional on-device classifiers for memory need and moment importance, enabling selective recall and local summarization. Finally, we plan to investigate human-facing aspects further including social acceptability, accessibility, and personalization, supported by open-source release of our datasets and design toolkit for community replication.

### v2

@@ 3.6. Training and Profiling
To design a model that fits the constraints of our ear-mounted device, training and profiling was done using a commercial TinyML platform, Edge Impulse, targeting reference microcontrollers comparable to the XIAO ESP32-S3 Sense used in our prototype. The wake word detector operates on 1-second windows of 16 kHz mono audio with a 500 ms stride, producing overlapping predictions suitable for continuous listening.
@@
The quantized model achieved 73.9 % validation accuracy (AUC = 0.92), with the strongest separation between the wake word and background noise. Remaining confusions occurred mainly between “hey dotty” and near-rhymes such as “hey dobby,” which we treated conservatively as non-triggers. Despite the synthetic nature of the dataset, this performance was sufficient for real-time triggering on-device.
@@ 4. DISCUSSION
Responsiveness on a constrained form factor (RQ1). Achieving interactive latency within a small ear-mounted device required restricting the wearable to wake word detection, capture, and streaming while offloading all inference to a paired smartphone. This allowed us to keep the device lightweight and thermally safe. Despite limited battery capacity and compute, we achieved query latencies of 2–3 s, showing that fully local multimodal interaction is feasible without cloud support.
Balancing energy, latency, and bandwidth (RQ2). Our design adopts a hybrid sensing pipeline: BLE for provisioning, Wi-Fi for runtime streaming, and short-lived event-driven sessions to reduce idle drain. Audio is compressed via 4-bit ADPCM, and single-frame snapshots replace continuous video. These decisions reduced power and network load while maintaining interactive performance.
Designing for privacy and acceptability (RQ3). Three properties shaped trust: (i) interaction is explicitly triggered by the user via wake word, (ii) all sensor data is processed locally, with no cloud calls, and (iii) the form factor is subtle and familiar. This design made the device feel unobtrusive in public. Still, some ambiguity remained; future iterations may include subtle LED or haptic cues to clarify capture status for both wearer and bystanders.
@@
System co-design. Many integration challenges (e.g., Wi-Fi/BLE coexistence, buffer stability, antenna layout) only emerged late in iteration. Prototyping tools like the early webserver example allowed us to validate hardware, streaming paths, and thermal behavior before inference was integrated.
@@ 4.1. Limitations
Despite these advances, several limitations remain. The evaluation involved a small number of short-duration deployments and relied primarily on internal testing. Longer, in-the-wild studies are necessary to assess reliability, comfort, and social perception over time. Our analysis also focused on a single smartphone class (iPhone 14); behavior on lower-end hardware or across operating systems may differ. Finally, while the current prototype supports wake word activation and multimodal inference, it does not yet integrate higher-level capabilities such as contextual memory or continuous dialogue—areas we plan to address next.
@@ 4.2. Future Work
We are in the process of obtaining IRB approval for a pilot study to evaluate everyday usability. We plan to collect both quantitative and qualitative data, including end-to-end latency, wake word accuracy per-participant, Wi-Fi reconnection behavior, battery drop over sessions, and participant feedback on comfort, perceived privacy, and bystander reactions. Future work will expand both the scope and depth of the system. On the hardware and networking side, we plan extended deployments, adaptive power management, and modular enclosures to support additional sensors and continuous use. Alternatives to wake word activation will also be explored further, including integration of a tactile sensor. On the intelligence layer, we plan to complete development of additional on-device classifiers for memory need and moment importance, enabling selective recall and local summarization. Finally, we plan to investigate human-facing aspects including social acceptability, accessibility, and personalization, supported by open-source release of our datasets and design toolkit for community replication.